# Scaling Environments for Organoid Intelligence with LLM-Automated Design and Plasticity-Based Evaluation

### v1

@@ Abstract
As the complexity of artificial agents increases, the design of environments that can effectively shape their behavior and capabilities has become a critical research frontier. We propose a framework that extends this principle to a novel class of agents: biological neural networks in the form of neural organoids. This paper introduces three scalable, closed-loop virtual environments designed to train organoid-based biological agents and probe the underlying mechanisms of learning, such as long-term potentiation (LTP) and long-term depression (LTD). We detail the design of three distinct task environments with increasing complexity: (1) a conditional avoidance task, (2) a one-dimensional predator-prey scenario, and (3) a replication of the classic Pong game. For each environment, we formalize the state and action spaces, the sensory encoding and motor decoding mechanisms, and the feedback protocols based on predictable (reward) and unpredictable (punishment) stimulation. Furthermore, we propose a novel meta-learning approach where a Large Language Model (LLM) is used to automate the generation and optimization of experimental protocols, scaling the process of environment and curriculum design. Finally, we outline a multi-modal approach for evaluating learning by measuring synaptic plasticity at electrophysiological, cellular, and molecular levels. This work bridges the gap between computational neuroscience and agent-based AI, offering a unique platform for studying embodiment, learning, and intelligence in a controlled biological substrate.
@@ 1 Introduction
The parallel scaling of model size, dataset size, and computation has yielded remarkable emergent capabilities in Large Language Models (LLMs) (Brown et al., 2020). A crucial, yet comparatively underexplored, dimension for advancing agent intelligence is the scaling of environments. Environments provide the essential data for agents to acquire adaptive behaviors through interaction, moving beyond static imitation learning towards true end-to-end autonomy (Sutton and Barto, 2018). The richness, diversity, and compositional structure of an environment directly shape an agent’s ability to reason, plan, and generalize.
This paper explores these principles in the context of a novel agent substrate: living neural organoids. Recent advances in organoid technology (Lancaster et al., 2013) and brain-computer interfaces have enabled the creation of Organoid Intelligence (OI) systems, where 3D neural cultures are embodied within simulated environments (Smirnova et al., 2023; Kagan et al., 2022). These biological agents offer a unique opportunity to study the physical mechanisms of learning and memory, such as synaptic plasticity, in a human-derived neural system.
Here, we present a framework for designing, scaling, and evaluating simplified virtual environments for training organoid-based agents. We contribute three distinct experimental designs that form a curriculum of increasing complexity, focusing on fundamental agent capabilities like state avoidance, goal seeking, and dynamic interception. We formalize the environment infrastructure, including task formulation, action-space design, and agent integration via closed-loop electrophysiological stimulation and recording. Crucially, we also propose leveraging an LLM to automate the design of these experimental protocols, enabling a new paradigm for scaling research in this domain. By grounding agent learning in measurable biological changes like Long-Term Potentiation (LTP) and Long-Term Depression (LTD), our work offers a novel perspective on agent evaluation, moving beyond simple task performance to the underlying adaptation of the agent’s neural architecture.
@@ 2 Related work
The concept of interfacing living neural tissue with artificial environments to study learning has a rich history. Early work with dissociated 2D neuronal cultures on multi-electrode arrays (MEAs) established the principle of in vitro embodiment, where neurally controlled animats learned to navigate a simulated world within a closed-loop system (DeMarse et al., 2001). This foundational research demonstrated that biological neural networks could exhibit goal-directed behavior when provided with structured sensory feedback and the means to act upon their environment.
The transition from 2D cultures to 3D neural organoids represents a significant increase in biological fidelity, offering a model that more closely recapitulates the architecture, cellular diversity, and developmental programs of the human brain (Lancaster et al., 2013). This complexity makes them a more powerful substrate for investigating network-level phenomena relevant to agent learning.
The free-engery principle (Friston, 2010) shows that predictable stimuli function as a reward (reinforcing actions that made the environment more predictable), while unpredictable, high-entropy stimuli act as a punishment. This demonstrated that goal-directed learning could be induced by structuring sensory feedback to align with the network’s intrinsic drive to minimize surprise.
Furthermore, we incorporate a novel methodological approach by using an LLM to automate the design of experimental protocols. This aligns with an emerging trend of using AI for autonomous scientific discovery, where models have successfully optimized complex tasks in chemistry and materials science (Boiko et al., 2023), suggesting their immense potential to accelerate and scale research in computational neuroscience and agent training.
@@ 3 Methods
3.1 Multi-electrode array
All proposed experiments leverage a multi-electrode array (MEA) platform, capable of concurrent stimulation and recording from a biological culture. We assume a setup with four distinct electrode groups (A, B, C, D), making this setup simple to recreate. Each electrode group can contain multiple electrodes. Groups A and B are designated for recording neural activity (motor output), while groups C and D are used for stimulation (sensory input).
3.2 Biological substrates
@@
While this paper focuses on neural organoids due to their structural complexity mimicking the human brain, the proposed experimental framework is fundamentally substrate-agnostic. These environments can be readily adapted for other biological preparations, including neural spheroids, 2D neuronal cultures, brain-on-a-chip platforms, or even non-neural tissue. The core requirement is a biological system that exhibits plasticity and can be integrated into a closed-loop system via an MEA for concurrent stimulation and recording. This flexibility allows for comparative studies across different levels of biological organization and complexity.
3.3 Agent feedback: Implementing reinforcement via the free-energy principle
@@
The learning framework is predicated on principles of reinforcement learning (RL), where agent behavior is shaped by feedback from the environment. We operationalize this feedback through the lens of the free-energy principle (Friston, 2010), which posits that biological systems act to minimize surprise or prediction error. In this context, feedback signals are designed to be either predictable (low surprise) or unpredictable (high surprise), serving as proxies for positive and negative reinforcement, respectively. This mechanism provides the necessary error signals to drive synaptic plasticity, the biological substrate of learning.
@@
Reward (positive reinforcement). A reward signal, corresponding to a positive outcome (e.g., capturing prey, intercepting the ball), is delivered as a predictable, low-entropy electrical stimulus to every electrode. A concrete implementation involves delivering a consistent, low-frequency sine wave across all stimulation electrodes. This predictable sensory input minimizes surprise and serves as a positive reward. Alternatively, a dose of dopamine can be delivered to the agent via UV light-controlled uncaging, which plays a role in inducing long-term potentiation (Schultz et al., 1997).
@@
Punishment (negative reinforcement). Conversely, a punishment signal, corresponding to a negative outcome (e.g., entering an aversive zone, missing the ball), is delivered as an unpredictable, high-entropy stimulus. This is implemented by applying a white-noise electrical signal, characterized by random amplitude and frequency, to a randomly selected subset of stimulation electrodes. This chaotic and surprising sensory input functions as an aversive signal or punishment, thereby acting as a powerful error signal. The objective is to drive the network to alter its policy to avoid states and actions that lead to such high-surprise outcomes.
@@
Together, this bipolar feedback scheme allows the agent to construct an internal model of its environment. By selectively reinforcing successful action sequences with predictable rewards and suppressing unsuccessful ones with unpredictable punishments, the framework directly translates the abstract principles of reinforcement learning into tangible biophysical changes within the neural network.
Encoding state. Information about the world state, such as the location of the organoid, can be encoded by stimulating a specific electrode. If there are 8 electrodes in a group, each can correspond to a different location. If there is only 1 electrode available, the location can be encoded by the frequency of the stimulation.
3.4 Benchmarks and evaluation: Measuring learning as plasticity
@@
A unique advantage of using a biological agent is the ability to move beyond purely behavioral metrics and directly measure learning as a physical reorganization of the neural substrate. Our framework proposes a multi-scale evaluation strategy to quantify synaptic plasticity, correlating improvements in task performance with underlying biological changes. At the functional level, we can probe network connectivity and synaptic efficacy by measuring field excitatory postsynaptic potentials (fEPSPs) before and after training epochs. By delivering a test pulse to a stimulation electrode and recording the evoked potential in a downstream population, we can measure the fEPSP slope as a proxy for synaptic strength. A sustained increase in this slope post-training serves as a canonical indicator of long-term potentiation (LTP), while a decrease signifies long-term depression (LTD), providing a direct link between the agent’s experience and Hebbian plasticity (Bliss and Lomo, 1973; Malenka and Bear, 2004).
@@
To gain insight into how neural populations encode task variables, we can employ optical imaging at cellular resolution. For organoids expressing Genetically Encoded Calcium Indicators (GECIs) like GCaMP (Chen et al., 2013), two-photon microscopy allows for longitudinal tracking of the activity of hundreds of individual neurons. This approach enables us to observe how network dynamics adapt over the course of learning, identifying the emergence of stable neural ensembles that may represent specific game states, actions, or environmental features. Changes in the synchrony, frequency, and spatial patterns of these calcium transients can elucidate how the agent’s internal model of the environment is formed and refined through interaction.
@@
Finally, at the conclusion of an experiment, we can perform post-hoc molecular analyses to visualize the structural and chemical correlates of the observed functional changes. Immunohistochemistry can be used to probe the molecular machinery of plasticity. For instance, staining for the trafficking of AMPA and NMDA receptor subunits can reveal changes in synaptic receptor density, a key mechanism underlying the expression of LTP and LTD (Malenka and Bear, 2004). Furthermore, using antibodies against phosphorylated proteins that are critical for plasticity, such as pCaMKII or pGluA1, provides a molecular snapshot of synapses that were recently potentiated, thereby linking task-relevant activity to specific structural changes (Heo et al., 2023). By integrating these electrophysiological, optical, and molecular benchmarks, we can construct a comprehensive picture of learning that bridges behavioral performance with the fundamental principles of neural adaptation.
@@ 4.1 Environment 1: Conditional avoidance
This initial experiment models a classic aversive conditioning task where the biological agent learns to associate a spatial region with negative stimuli and actively avoid it. This environment is inspired by similiar psychological experiments on mice (Miller, 1948). A schematic of the environment is shown in Figure 1.
@@ 4.1.1 Protocol description
    State Representation and Sensory Encoding: The agent’s location within an 8-position one-dimensional grid is conveyed via patterned electrical stimulation. Each position (1-8) maps to a unique spatio-temporal stimulation pattern delivered across electrode groups C and D. This provides the organoid with proprioceptive information about its virtual embodiment. Positions 1-5 constitute a safe zone, while positions 6-8 form an aversive zone.
@@
    Action Decoding and Execution: The agent’s intended movement is decoded by comparing the aggregate spike counts recorded from groups A and B over a discrete time window TT. If the spike count from group A exceeds that of group B, the agent’s virtual position is decremented (move left). Conversely, if group B’s activity is higher, the position is incremented (move right). This mechanism translates differential network activity into a binary action space.
@@
    Feedback and State Transition: The environment provides feedback based on the agent’s position. If the agent enters the aversive zone (positions 6-8), it receives an unpredictable, high-entropy electrical stimulus (punishment). The intensity of this stimulus is scaled with the depth of incursion (position 8>7>68>7>6), providing a gradient for learning. Conversely, remaining in the safe zone for a sustained period of ZZ timesteps triggers a predictable reward.
4.1.2 Environmental scaling

### v2

# Published in the proceedings of the 39th Conference on Neural Information Processing Systems (NeurIPS 2025) Workshop: Scaling Environments for Agents (SEA). Additionally accepted for presentation in NeurIPS 2025 Workshop: Embodied World Models for Decision Making

@@ Abstract
The capacity of an embodied agent to understand, predict, and interact with its environment is fundamentally contingent on an internal world model. This paper introduces a novel framework for investigating the formation and adaptation of such world models within a biological substrate: human neural organoids. We present a curriculum of three scalable, closed-loop virtual environments designed to train these biological agents and probe the underlying synaptic mechanisms of learning, such as long-term potentiation (LTP) and long-term depression (LTD). We detail the design of three distinct task environments that demand progressively more sophisticated world models for successful decision-making: (1) a conditional avoidance task for learning static state-action contingencies, (2) a one-dimensional predator-prey scenario for goal-directed interaction, and (3) a replication of the classic Pong game for modeling dynamic, continuous-time systems. For each environment, we formalize the state and action spaces, the sensory encoding and motor decoding mechanisms, and the feedback protocols based on predictable (reward) and unpredictable (punishment) stimulation, which serve to drive model refinement. In a significant methodological advance, we propose a meta-learning approach where a Large Language Model (LLM) automates the generative design and optimization of experimental protocols, thereby scaling the process of environment and curriculum design. Finally, we outline a multi-modal evaluation strategy that moves beyond task performance to directly measure the physical correlates of the learned world model by quantifying synaptic plasticity at electrophysiological, cellular, and molecular levels. This work bridges the gap between model-based reinforcement learning and computational neuroscience, offering a unique platform for studying embodiment, decision-making, and the physical basis of intelligence.
@@ 1 Introduction
The development of sophisticated world models has emerged as a cornerstone of modern embodied artificial intelligence, enabling agents to reason, plan, and act effectively in complex environments (Sutton and Barto, 2018). While progress in AI has been propelled by the concurrent scaling of model architectures, datasets, and computational resources (Brown et al., 2020), the crucial role of the environment itself as a driver of intelligence remains a frontier for exploration. The structure, complexity, and interactivity of an agent’s world directly constrain and shape the internal models it can acquire. Rich, dynamic environments provide the necessary data streams for agents to move beyond passive prediction and static imitation, fostering the development of robust, goal-directed behaviors through continuous interaction.
This paper explores these principles through the lens of a novel and powerful agent substrate: living neural organoids. Recent breakthroughs in stem cell biology (Lancaster et al., 2013) and neuro-interfacing technologies have converged to create the field of Organoid Intelligence (OI), where three-dimensional human neural cultures are embodied within simulated, interactive worlds (Smirnova et al., 2023; Kagan et al., 2022). These biological agents, derived from human cells, present an unprecedented opportunity to study the formation of learned world models not as an abstract computational process, but as a tangible biological phenomenon rooted in the physical principles of synaptic plasticity.
Here, we present a comprehensive framework for designing, scaling, and evaluating virtual environments specifically tailored to induce and investigate learning in organoid-based agents. We make three primary contributions. First, we introduce a curriculum of three distinct environments, each designed to probe a different facet of decision-making and world modeling, ranging from simple state avoidance to dynamic, long-horizon interception tasks. We formalize the complete closed-loop system, including sensory state encoding, motor action decoding, and a reinforcement mechanism grounded in neuroscientific principles. Second, we propose a novel approach for evaluating an agent’s learned world model, moving beyond behavioral metrics to directly measure the underlying biophysical changes in synaptic strength, namely Long-Term Potentiation (LTP) and Long-Term Depression (LTD). This allows us to correlate task proficiency with the physical reorganization of the agent’s neural architecture. Third, we introduce a generative, meta-learning framework that leverages a Large Language Model (LLM) to automate the design and optimization of experimental protocols. This positions the LLM as a tool for goal-directed environment generation, enabling a new paradigm for scaling and accelerating research into embodied intelligence. By wedding model-based reinforcement learning concepts with a physical biological system, our work offers a unique platform to advance our understanding of how embodied agents learn to understand, predict, and interact with their world.
@@ 2 Related work
The concept of embodying living neural tissue within artificial environments to study learning and adaptation has a rich history. Foundational work with dissociated 2D neuronal cultures on multi-electrode arrays (MEAs) established the principle of in vitro embodiment, demonstrating that neurally-controlled animats could learn to navigate simulated worlds and exhibit goal-directed behavior when provided with a closed-loop system of sensory feedback and motor control (DeMarse et al., 2001). This research provided the first proof-of-concept that biological neural networks, de-coupled from a biological body, could form instrumental relationships with an external, virtual world.
The evolution from flat, 2D cultures to 3D neural organoids marks a profound leap in biological fidelity. Organoids more closely recapitulate the complex cytoarchitecture, cellular diversity, and developmental trajectories of the human brain, offering a far richer substrate for investigating the network-level phenomena that underpin complex decision-making and the formation of internal models (Lancaster et al., 2013). Recent work has successfully demonstrated that these organoid-based systems can learn to perform tasks, such as playing a simplified version of the game Pong, by harnessing their inherent plasticity (Kagan et al., 2022).
Our framework builds directly upon these advancements by grounding the learning mechanism in the free-energy principle (Friston, 2010). This theoretical framework posits that self-organizing systems, including the brain, act to minimize prediction error or surprise. Within this paradigm, predictable sensory stimuli can be interpreted as an intrinsic reward, reinforcing the policies that led to them, while unpredictable, high-entropy stimuli act as an aversive signal or punishment, indicating a failure of the agent’s internal world model. This provides a powerful, neuro-plausible mechanism for implementing model-based reinforcement learning, where feedback directly drives the refinement of the agent’s predictive model of its environment.
Furthermore, our work introduces a significant methodological innovation by employing an LLM to automate the design of experimental protocols. This aligns with an emergent trend of leveraging AI for autonomous scientific discovery. Such systems have already demonstrated success in optimizing complex processes in domains like chemistry and materials science (Boiko et al., 2023). By applying this paradigm to the training of embodied agents, we propose a system where the LLM acts as a meta-controller, designing the very worlds and curricula in which the biological agent learns. This approach addresses leveraging the knowledge within LLMs to guide agent decision-making, extending the concept to guide the entire learning and discovery process itself.
@@ 3 Methods
3.1 Multi-electrode array interface
All proposed experiments are built upon a multi-electrode array (MEA) platform, which provides the critical bidirectional interface for embodying the biological agent. This platform must be capable of concurrent, high-fidelity stimulation and recording from the neural culture. For simplicity and reproducibility, we assume a setup with at least four distinct electrode groups (A, B, C, D), which can be spatially organized across the array. Each group can consist of one or more electrodes. Groups A and B are designated for recording spontaneous and evoked neural activity, serving as the agent’s motor output channels. Groups C and D are dedicated to delivering patterned electrical stimuli, functioning as the agent’s sensory input channels. This architecture forms the physical link between the virtual environment and the biological substrate.
3.2 Biological substrates and embodiment
@@
While this paper focuses on human neural organoids due to their advanced structural and cellular complexity that emulates aspects of the developing human brain, the experimental framework we propose is fundamentally substrate-agnostic. The core principles of sensory encoding, action decoding, and feedback-driven learning can be readily adapted to a variety of biological preparations. These include simpler neural spheroids, traditional 2D neuronal cultures, more structured brain-on-a-chip platforms, or even non-neural tissues that exhibit adaptive plasticity. The essential requirement is a biological system that can undergo experience-dependent changes and can be integrated into a closed-loop control system via an MEA for real-time stimulation and recording. This inherent flexibility allows for powerful comparative studies across different scales of biological organization, from dissociated networks to complex, multi-layered tissues.
3.3 Inducing world model formation via predictive coding and reinforcement
@@
The core learning mechanism of our framework is predicated on the principles of model-based reinforcement learning (RL), operationalized through the neuro-plausible lens of the free-energy principle (Friston, 2010). This principle suggests that biological agents inherently act to minimize the discrepancy between their predictions about the world and the sensory evidence they receive, a quantity known as surprise or prediction error. In our framework, we leverage this concept by designing feedback signals to be either highly predictable (low surprise) or highly unpredictable (high surprise), serving as functional proxies for positive and negative reinforcement, respectively. This mechanism provides the essential error signals required to drive synaptic plasticity, the biological substrate for updating the agent’s internal world model.
@@
Reward (Model Confirmation). A reward signal, delivered upon achieving a desirable state (e.g., capturing prey), is implemented as a predictable, low-entropy electrical stimulus. A concrete implementation involves delivering a consistent, low-frequency sinusoidal wave across all stimulation electrodes. This spatially and temporally predictable sensory input minimizes surprise, serving as a confirmation signal that reinforces the agent’s preceding policy and strengthens its internal model of successful state-action sequences. An alternative biological implementation involves the targeted delivery of dopamine via UV light-controlled uncaging, directly activating the brain’s reward pathways known to be crucial for inducing long-term potentiation (Schultz et al., 1997).
@@
Punishment (Model Violation). Conversely, a punishment signal, corresponding to an undesirable outcome (e.g., entering an aversive zone), is delivered as an unpredictable, high-entropy stimulus. This is implemented by applying a white-noise electrical signal, characterized by random amplitude and frequency, to a randomly selected subset of stimulation electrodes. This chaotic and surprising sensory input functions as a powerful error signal, indicating a violation of the agent’s predictive model. The objective is to drive the network to update its policy to avoid states and actions that lead to such high-surprise, model-violating outcomes.
@@
This bipolar feedback scheme directly translates the abstract principles of reinforcement learning into tangible biophysical drivers of change. By selectively confirming successful policies with predictable rewards and signaling model failures with unpredictable punishments, the framework compels the biological neural network to iteratively build and refine an internal model that can better predict and control its virtual environment.
Encoding State. To enable the formation of a world model, the agent must receive information about the state of its environment. The agent’s location, or the location of objects in the virtual world, is encoded through patterned electrical stimulation. For a discrete state space with 8 locations, each location can be mapped to a specific stimulation electrode within a sensory group. For continuous or higher-dimensional states, information can be rate-coded, where the frequency of stimulation on a given electrode corresponds to a variable like distance or position along an axis. This provides the raw sensory data from which the agent must construct its model.
3.4 Evaluating the embodied world model: From task performance to synaptic correlates
A profound advantage of employing a biological agent is the ability to move beyond simplistic behavioral metrics and directly measure learning as the physical reorganization of the agent’s underlying neural substrate. Our framework proposes a multi-scale evaluation strategy to quantify this plasticity, allowing us to correlate improvements in decision-making with the specific biological changes that constitute the learned world model.
@@ 3.4 Evaluating the embodied world model: From task performance to synaptic correlates
At the functional network level, we can probe synaptic efficacy by measuring field excitatory postsynaptic potentials (fEPSPs) before and after training epochs. By delivering a controlled test pulse to a stimulation electrode and recording the evoked potential in a downstream population, the slope of the fEPSP serves as a robust proxy for synaptic strength. A sustained increase in this slope post-training is the canonical electrophysiological signature of long-term potentiation (LTP), while a decrease signifies long-term depression (LTD). This provides a direct, quantifiable link between the agent’s interactive experience and the Hebbian plasticity mechanisms thought to underlie learning and memory (Bliss and Lomo, 1973; Malenka and Bear, 2004).
@@
To resolve how neural populations encode task variables and represent the world model at a finer scale, we can employ optical imaging. In organoids engineered to express Genetically Encoded Calcium Indicators (GECIs) like GCaMP (Chen et al., 2013), two-photon microscopy enables longitudinal tracking of the activity of hundreds of individual neurons. This technique allows us to observe the evolution of network dynamics as learning progresses, potentially identifying the emergence of stable neural ensembles that represent specific states, actions, or environmental features. Changes in the synchrony, firing rate, and spatial organization of these calcium transients can elucidate how the agent’s internal model is formed, stored, and refined through goal-directed interaction.
@@
Finally, at the conclusion of an experiment, post-hoc molecular analyses can reveal the structural correlates of the observed functional changes. Immunohistochemistry can be used to visualize the molecular machinery of plasticity. For example, staining for the differential trafficking of AMPA and NMDA receptor subunits can expose changes in synaptic receptor density, a key mechanism underlying the expression of LTP and LTD (Malenka and Bear, 2004). Furthermore, using antibodies against plasticity-related phosphorylated proteins, such as pCaMKII or pGluA1, provides a molecular snapshot of synapses that were recently potentiated, thereby linking task-relevant neural activity to specific structural modifications (Heo et al., 2023). By integrating these electrophysiological, optical, and molecular benchmarks, we can construct a comprehensive portrait of learning that bridges behavioral performance with the fundamental principles of neural adaptation, offering a ground-truth look into the learned world model.
@@ 4 Environments
We propose a curriculum of three virtual environments, each designed to incrementally increase the demands on the agent’s world modeling and decision-making capabilities.
@@ 4.1 Environment 1: Conditional avoidance
This foundational experiment implements a classic aversive conditioning task, designed to test the agent’s ability to build a simple world model that associates a specific region of its state space with negative outcomes and to develop a policy to actively avoid it. This environment, inspired by foundational experiments in learning theory (Miller, 1948), probes the most basic form of goal-directed behavior: avoidance of harm. A schematic is shown in Figure 1.
@@ 4.1.1 Protocol description
    State Representation and Sensory Encoding: The agent is embodied as a point in a one-dimensional grid of 8 positions. Its current location is continuously conveyed via a unique spatio-temporal electrical stimulation pattern delivered across sensory electrode groups C and D. This provides the organoid with proprioceptive information about its virtual location. The state space is partitioned into a safe zone (positions 1-5) and an aversive zone (positions 6-8).
@@
    Action Decoding and Execution: The agent’s intended movement is decoded by comparing the aggregate spike counts recorded from motor output groups A and B over a discrete time window TT. If the integrated activity from group A exceeds that of group B, the agent’s virtual position is decremented (action: move left). Conversely, if group B’s activity is dominant, the position is incremented (action: move right). This establishes a simple, differential mapping from network activity to a binary action space.
@@
    Feedback and State Transition: The environment delivers feedback contingent on the agent’s state. Upon entering the aversive zone, the agent receives an unpredictable, high-entropy electrical stimulus (punishment). The intensity of this stimulus is scaled with the depth of incursion (e.g., amplitude at position 8>7>68>7>6), providing a gradient that can accelerate learning. Conversely, successfully remaining in the safe zone for a sustained duration of ZZ timesteps triggers a predictable, low-entropy stimulus (reward).
4.1.2 Environmental scaling and long-horizon planning